An equipment costing 5, 000, 000 can be sold for 1, 000, 000 after it’s useful life of five years. What will be the book value of the said equipment if the company plans to sell it the end of 3 years? Use the sum of years digit method.
Found 2 solutions by ikleyn, Thing_a_majig:
Answer by ikleyn(52781)

.
An equipment costing 5, 000, 000 can be sold for 1, 000, 000 after it’s useful life of five years.
What will be the book value of the said equipment if the company plans to sell it the end of 3 years?
Use the sum of years digit method.
~~~~~~~~~~~~~~~~


Is the term  " sum of years digit method "  your own invention ?


If  " YES ",  did you patent or,  at least,  published it ?



Do it urgently . . .


/////////////////


What I wrote above,  were jokes.

What is written below,  is absolutely serious.


        - - - NEVER  print such  GIBBERISH  in the future. - - -




Answer by Thing_a_majig(1)

If the useful life of the equipment is 5 years, the equipment would have depreciated by 5,000,000 - 1,000,000 = 4,000,000 in currency within that period.

To do sum-of-years digits depreciation, add the ‘digits’ in its useful life: 1 + 2 + 3 + 4 + 5 = 15.

The amount of annual depreciation is then computed as follows --

1st year: (5/15)×4,000,000=1,333,333.33
2nd year: (4/15)×4,000,000=1,066,666.67
3rd year: (3/15)×4,000,000=800,000
4th year: (2/15)×4,000,000=533,333.33
5th year: (1/15)×4,000,000=266,666.67

Therefore the book value of the equipment after 3 years is 5,000,000 - (1,333,333.33+ 1,066,666.67+ 800,000)= 1,800,000.

***The sum-of-years digits method is a standard procedure in depreciating the value of an asset,
and most of the depreciation associated with it is recognized in the first few years of its useful life.